If you divide or multiply each number in a pythagorean triple by a constant will you get another pythagorean triple? Is this statement true or false, if false, why?
1
Expert's answer
2012-11-05T09:57:03-0500
As we know c^2=a^2+b^2. As we multiply all three numbers bysome d we get:
(dc)^2=(ad)^2+(db)^2 ord^2c^2=d^2(a^2+b^2) which is right. So pythagorean equality is still working we
got another pythagorean triple.